# references.bib

### v1

@@ Appendix F Discussion
  • In this work, we simply collect all studies from our health system, resulting in an imbalanced pre-training dataset. Building on findings from the natural image domain, we believe that developing a systematic approach to construct a relatively balanced pre-training dataset is an important direction for future work.
Broader Impacts. HLIP models pre-trained on our collected datasets, BrainMRI220K and HeadCT240K, will be released shortly. Since HLIP can make predictions directly on raw radiology studies, it offers flexibility for real-world clinical use and may serve as a reference tool for radiologists, potentially alleviating their workload. However, despite HLIP’s strong performance on several benchmarks, it should not be solely relied upon for diagnostic decisions.

### v2

@@ Appendix E Discussion
Study Curation. In this work, we collect all studies from our health system, resulting in an imbalanced pre-training dataset. Building on findings from the natural image domain, we believe that developing a systematic approach to construct a more balanced pre-training dataset is an important direction for future work. Please note that this curation should occur at the study level and therefore does not impose additional burden on radiologists, unlike selecting representative scans or slices.
Moreover, we observe that zero-shot transferability does not always correlate with the number of studies used for training. For example, although keyword search yields more patients with meningioma or metastasis than with glioma, zero-shot performance on glioma is substantially higher. While tumor size may partially explain this discrepancy, we believe it warrants further investigation in future work.
Scale. Even at its current scale of 200K studies, the largest to date, our dataset remains far smaller than training scales in the natural image domain, which typically involve billions of samples. Although collected from a neuro-radiology system, we find that other organ systems such as the spine, abdomen, and knee occasionally appear in our dataset, and the current model demonstrates promising zero-shot transferability to these organs as well. Therefore, we believe there is no barrier to training a model on more comprehensive data, including different organ systems (e.g., brain, cardiac, chest, abdomen) and modalities (e.g., CT and MRI). Since no human labeling burden is involved, we expect the training scale to expand to millions of samples in the near future.
As the dataset scales further, especially when encompassing multiple organ systems and modalities, the underlying patterns will become more complex. Therefore, we believe that larger batch sizes will be necessary. Although our hierarchical attention mechanism introduces no additional computational burden, we rely on both flash attention and gradient checkpointing to achieve a batch size of 256 on 8 L40 GPUs. Further increasing the batch size will require either a gradient accumulation strategy or additional computational resources.
Moreover, as the dataset scales up, larger models may prove beneficial but will also require additional computational resources.
Vision-Language Model. Although current vision–language models achieve impressive performance even on medical tasks, we believe these general-purpose models will reach a bottleneck when sufficient specialist medical data are not available on the internet. On the other hand, as with other language-supervised visual encoders, HLIP’s visual encoder could facilitate the development of vision language models under current frameworks. Therefore, we expect to see a radiology specialist vision–language model that can operate on entire studies in the near future.